OpenStudy (anonymous):

Find the sum of the exterior angle of a polygon angles of a polygon with 57 sides.

OpenStudy (campbell_st):

the sum of the exterior angles is always 360 degrees so in your question 1 exterior angle = 360/57 that is provided the polygon is regular

OpenStudy (anonymous):

but how can i get the sum of the exterior angle? so i would have to divide 360/57 ?

OpenStudy (campbell_st):

well start with the interior sum \[Angle Sum = (n - 2) \times 180\] so substitute n = 57 to find the angle sum . then divide the angle sum by 57 to find the size of 1 angle... then the 1 exterior angle will be supplementary.  so exterior angle = 180 - interior and there will be 57 of them so exterior x 57 for their sum.

OpenStudy (anonymous):

okay so 57-2x180 = 9900

OpenStudy (campbell_st):

yep so 1 interior angle = 9900/57 then 1 exterior = 180 = 9900/57 then answer x 57 and you'll get 360 leave your answer for 1 angle and the exterior angle as a mixed number.... don't use decimals and round

OpenStudy (anonymous):

okay so it would be 360 ? c:

OpenStudy (campbell_st):

exactly

OpenStudy (anonymous):

yay ! thank you c: